what is the vapour pressure of liquid gas solutions and how does raoults law become a special case of henrys law?

what is the  vapour pressure of liquid gas solutions and how does raoults law become a special case of henrys law?


 

Grade:

1 Answers

ruchi yadav
10 years ago

For a dilute solution, the concentration of the solute is approximately proportional to its mole fractionx, and Henry's law can be written as:

p = k_{\rm H}\,x

This can be compared withRaoult's Law:

p = p^\star\,x

wherep* is the vapor pressure of the pure component.

At first sight, Raoult's law appears to be a special case of Henry's law wherekH=p*. This is true for pairs of closely related substances, such asbenzeneandtoluene, which obey Raoult's law over the entire composition range: such mixtures are called "ideal mixtures".

The general case is that both laws arelimit laws, and they apply at opposite ends of the composition range. The vapor pressure of the component in large excess, such as the solvent for a dilute solution, is proportional to its mole fraction, and the constant of proportionality is the vapor pressure of the pure substance (Raoult's law). The vapor pressure of the solute is also proportional to the solute's mole fraction, but the constant of proportionality is different and must be determined experimentally (Henry's law). In mathematical terms:

Raoult's law:\lim_{x\to 1}\left( \frac{p}{x}\right) = p^\star
Henry's law:\lim_{x\to 0}\left( \frac{p}{x}\right) = k_{\rm H}

Raoult's law can also be related to non-gas solutes.



Thank You

Ruchi

Askiitians Faculty